Operating Leverage. You estimate that your cattle farm will generate $1 million of profits on sales of $4 million under normal economic conditions, and that the degree of operating leverage is 7.5. What will profits be if sales turn out to be $3.5 million? What if they are $4.5 million?
Answer #1

DOL =% change in profit/ % change in sales = 7.5

If sales = $3.5 m, % change in sales = (3.5-4)/4 =-12.5%

Hence % change in profit = 7.5*12.5% = -93.75%

Profit =1000000- 1000,000*93.75% = $62500

If sales = $4.5 m, % change in sales = (4.5- 4)/4.5 =11.11%

Hence % change in profit = 7.5*11.11% = 83.325%

Profit = 1000,000+ 1000000*83.325% = $1833250